# Data Guard: A Fine-grained Purpose-based Access Control System for Large Data Warehouses

### v1

@@ 5.1.1. Policy Compiler
We focus on the implementation of the $mask$ operator to perform column-level masking i.e. the data element to be masked is inside an attribute of $\mathcal{R}$. A naive approach to implementing the data-masking operator for a relation-valued attribute is to unnest the attribute, mask the data at a given field path within the attribute and reconstruct the attribute into its original structure. However, this approach produces complex SQL queries for relation-valued attributes, which can be both hard to read and debug especially when involving transformations on deeply nested attributes. We therefore implement a scalar version of the $mask$ operator for column-level masking. Specifically, the operator employs a user-defined function (UDF) MASK_FIELD(), which takes a field path $f$ as an argument and sets the value at the field path to $NULL$.
Consider a field path $f$ inside an attribute $\mathcal{A}$ of a relation $\mathcal{R}$. We start with an implementation of the column-masking UDF with the following signature:
MASK_FIELD($\mathcal{A}$: ANY, $f$: VARCHAR)
@@
The above UDF performs in-place update on attribute $\mathcal{A}$, instead of a sequence of unnest, transform, and re-construct operations. The return type of the UDF is the same as the data type of $\mathcal{A}$ to ensure the view schema is identical to the underlying table schema. The data-masking operator introduced in Equation 1 can thus be implemented by combining the MASK_FIELD() UDF with the CASE statement in SQL as follows:
CASE
WHEN p.cond THEN A
ELSE MASK_FIELD(A, f)
END
Now suppose we have two field paths $f_{1}$ and $f_{2}$, both accessing the same attribute $\mathcal{A}$ with matching policies $p_{1}$ and $p_{2}$ respectively. Algorithm 1 yields the following relational algebra plan:
$mask(mask(\mathcal{R},p_{1},f_{1}),p_{2},f_{2})$
which is translated into two consecutive projection operators. Since SQL engines commonly apply the projection merge optimization rule to combine consecutive projections in a plan into a single projection, the above plan is translated into the following SQL code:
CASE
WHEN p2.cond THEN (CASE WHEN p1.cond THEN A ELSE MASK_FIELD(A, f1) END)
ELSE MASK_FIELD((CASE WHEN p1.cond THEN A ELSE MASK_FIELD(A, f1) END) , f2)
END
In this expression, MASK_FIELD() is called once on $f_{2}$ and twice on $f_{1}$. A straightforward analysis shows that for $n$ fields accessing the same attribute $\mathcal{A}$, the number of UDF calls in the generated SQL will equal $12^{n}-$ 1.
We avoid the exponential number of UDF calls in the view SQL by making the policy condition a parameter of the UDF instead of using the CASE expression. This results in the following modified signature for the masking UDF:
MASK_FIELD_IF(cond: BOOLEAN, $\mathcal{A}$: ANY, $f$: VARCHAR) where the UDF only applies data masking when the condition is true and returns the original value of $\mathcal{A}$ otherwise.
With this optimization, the above nested CASE statement is rewritten with two UDF calls ($n$ UDF calls instead of $2^{n}$ for $n$ enforcements) as:
MASK_FIELD_IF(p2.cond,
MASK_FIELD_IF(p1.cond, A, f1), f2)
When the policy conditions are the same ($p_{1}.cond=p_{2}.cond$), the policy compiler further groups the nested UDF calls into a single UDF call:
MASK_FIELD_IF(p1.cond, A, CONCAT(f1, delim, f2))
5.1.2. Consent Bitmaps
@@
Recall from Section 4.1.3 that the most common type of data access policies in our system are consent-based, where the policy condition involves checking preferences of data subjects stored in a separate table, referred to as the lookup table. As described in Section 4.2, a naive implementation of policy compiler can produce views that join a table containing data subject information with the lookup table using the data subject id as the join key. At LinkedIn, this table stores the consents of more than 1 billion members and any joins against this table result in expensive data shuffles. In Spark for instance, such joins are implemented as Sort Merge Joins since the size of both the tables involved in the join operation are too large to be broadcast using Broadcast Hash Joins. Optimizing the view SQL for fetching attribute values from the lookup table is therefore critical to minimize the cost of adopting views.
@@
In order to address this scalability challenge, we pre-compute a consent bitmap for every consent attribute referenced in a policy definition and leverage statistics associated with each consent to minimize the bitmap size. Specifically, we use a ”true bitmap” (bitmap of data subject identifiers whose consent value is true) if minority of users have consent values set to true, and a false bitmap otherwise. Therefore, the number of user ids in each bitmap is at most half of the total number of users. Given that data subject identifiers at LinkedIn are numeric (a scenario we expect to be common), we use the Roaring Bitmap implementation  (Lemire et al., 2016) to compute the bitmaps as it achieves superior compression ratios compared to other compression methods. As an example, a Roaring bitmap of 200 million 8-byte integers that represent user ids has a size of around 134MB, which is small enough for the bitmaps to be loaded into memory on the worker nodes in Spark and Trino. In cases where identifiers are non-numeric, we can employ minimal perfect hashing functions (Belazzougui et al., 2009) to map keys into consecutive integers and then use Roaring Bitmap on the resulting integers.
@@
These bitmaps are then saved in a storage system like HDFS or cloud storage and are loaded at access time for fast in-memory lookup via a UDF. The bitmap implementation is much cheaper compared to a join-based solution, as we only introduce a small memory overhead for storing bitmaps in each worker’s memory. In order to abstract away the storage system implementation details, we wrap the bitmap loading and lookup actions into a BitmapManager interface, thus ensuring that the bitmap solution can be adopted across a range of different storage systems.
@@
To ensure freshness of bitmaps, a bitmap computation job is scheduled periodically to reflect the latest user consents in the generated snapshot. Each snapshot of a bitmap is addressed using a combination of bitmap name and a timestamp when the bitmap was generated. The different snapshots of a bitmap are stored in a time-partitioned folder. The bitmap lookup is wrapped inside a UDF with the following signature: HAS_USER_CONSENT(consents: VARCHAR, user_id: BIGINT, access_time: TIMESTAMP), where:
@@
  • access_time: the timestamp for determining the specific snapshot of a bitmap to load. This parameter gives users the ability to replay a query at a later time (aka time travel). By default, the data-masking views use the built-in SQL function ${CURRENT\_TIMESTAMP()}$ for the $access\_time$ parameter. The $access\_time$ parameter ensures that the same snapshot of bitmap is loaded across all the worker nodes for a given query.
5.1.3. Masking deduplication
Given the scale of LinkedIn’s data warehouse and the diversity of policies that need to be enforced, it is common to find tables in our warehouse containing attributes that are subject to multiple policies. Further, a single policy may apply to multiple attributes within the table. A naive implementation of the policy compiler in Algorithm 1 can result in redundant application of the $MASK\_FIELD\_IF()$ UDF.
@@ 5.1.3. Masking deduplication
  • If a nested field needs to be masked based on one or more policy conditions, each of which is applicable to an ancestor, then a masking operation on the nested field is unnecessary. For example, masking operation on the field path in row 2 of Table 1 is redundant given an identical policy condition applicable to the root path \$. Similarly, the masking on field path in row 4 is redundant given an identical policy condition applicable to its parent in row 3.
  • If a field in a table has two or more applicable policies and the set of consents used in policy $p_{1}$ is a superset of the set of consents used in a different policy $p_{i}$, then the masking with policy $p_{j}$ is unnecessary. For example, rows 5 and 6 of Table 1 show two different policies $p_{3}$ and $p_{4}$ applied to the same field path with overlapping consents. It is easy to see that policy condition check in row 6 is redundant and can be eliminated.
@@
1 \$currency-dollar\\$\$ $consent_{1}$ $p_{1}$
2 \$.$\$.col_{2}.field_{21}$ $consent_{1}$ $p_{1}$
3 \$.$\$.col_{3}$ $consent_{2}$ $p_{2}$
4 \$.$\$.col_{3}.[item].[?(@.field_{31}=^{\prime}s1^{\prime})]$ $consent_{2}$ $p_{2}$
5 \$.$\$.col_{4}.[value]$ $consent_{3}$ AND $consent_{4}$ $p_{3}$
6 \$.$\$.col_{4}.[value]$ $consent_{3}$ $p_{4}$
Table 1. Example of maskings with different field paths with applicable policies. Masking operations on field paths in rows 2, 4, and 6 can be pruned without impacting the masking results.
To eliminate such redundant masking operations, we construct a schema tree from a given list of field paths and their applicable policies. A schema tree is one where each node represents an operator of a field path expression defined in Definition 4.2 and nodes corresponding to successive operators of a field path expression have a parent-child relationship between them. A node that is a final operator of a field path has an attribute to store the list of all policies applicable to that field path. Figure 5(a) shows an example schema tree constructed from field paths in Table 1.
@@
Figure 5. An example schema tree constructed from field paths defined in Table 1.
After constructing the schema tree, we perform a pre-order traversal to prune redundant policies as detailed in Algorithm 2. The algorithm computes the minimal set of policies required to cover the set of consents of a node that do not overlap with any of its ancestors. This problem can be mapped to the well-known minimum set cover problem, which is an NP-complete problem (Garey and Johnson, 1990). We therefore use a greedy approximation algorithm that chooses policies with largest number of non-overlapping consents first. Figure 5(b) demonstrates the schema tree produced by applying Algorithm 2. Maskings in rows 1, 3, and 5 are chosen to be retained by the algorithm.
@@
         select $p\in P$ that maximizes $|p.consents\cap\Delta|$
         $\Delta\leftarrow\Delta\setminus p.consents$
@@
Suppose $col_{1}$ of $\mathcal{R}$ is labeled as a user id column, the resulting data-masking view SQL produced by Algorithm 1 is as follows:
@@ 5.1.4. View Maintenance
      If a view update is deemed necessary, the system invokes Algorithm 1 to generate a new view definition, and
@@
      registers the updated view definition with the view catalog (a Hive-based (Apache Hive, 2010) catalog in our ecosystem).
To ensure smooth rollout of views and enable rollbacks in case of errors, each data-masking view is versioned according to the semantic versioning scheme (Semantic Versioning, 2013). End user pipelines are routed to the latest view versions by default. We also provide users with the option to pin their pipelines to specific versions of views and be notified when newer versions of views become available. In this mode, users can test their pipelines against the new view versions before deploying the change to production. Older versions of views are garbage collected periodically to minimize resource consumption due to stale views.
5.2. View Consumption
At LinkedIn, warehouse users write data processing applications using Apache Spark and Trino as the underlying compute engines. To ensure ease-of-use of data-masking views in user applications, we implemented two important optimizations: (i) dynamic table-to-view routing, and (ii) multi-engine support, each of which will be discussed in detail in this section.
5.2.1. Dynamic Table-to-View Routing
A key design choice of our system was to ensure access control policies are automatically enforced during data accesses by redirecting table accesses to the appropriate data-masking views determined by the purpose of access. This minimizes overhead for existing warehouse users to ensure their accesses remain compliant. Recall from Section 3 that the table to view routing at access time is accomplished using a component called ViewShift. At LinkedIn, we implement this component as a plugin inside existing engine catalog implementations. ViewShift intercepts loadTable() API calls to the catalog from applications and returns the corresponding data-masking view instead. The ViewShift plugin exposes the following API:
@@ 5.2.1. Dynamic Table-to-View Routing
where $contextMap$ parameter is query context map that contains attributes such as the purpose of the access (obtained from the identity of the user that submits the query), the optional pinned version of the data masking view, and the environment where the application is running.
@@
5.2.2. Multi-engine support
In large organizations, it is common for infrastructure teams to support a diversity of data processing systems, each optimized for a specific set of use cases. For instance, Apache Spark is commonly used by data scientists for data exploration, while Trino is commonly used for ad hoc data analysis. There are variances in the SQL dialects supported by each system. Further, each system has a different mechanism and type system to define UDFs. A key requirement that Data Guard needed to address was to ensure that the data masking views are executable across different data processing systems and avoid view duplication by creating engine-specific views.
To ensure portability of views across engines, we leverage two in-house technologies, which are independently available as open-source projects: (i) a SQL translation tool called Coral (Coral, 2020) that translates SQL between different dialects, and (ii) a framework called Transport UDF (Transport, 2018) to author UDFs that can be executed on multiple engines.
@@ 5.2.2. Multi-engine support
The data-masking views generated by Data Guard are stored in a view catalog that supports the Hive SQL dialect and are translated by Coral to Spark/Trino SQL at access time. Coral uses Apache Calcite (Apache Calcite, 2014) under-the-hood to parse SQL from a source dialect into a relational algebra representation, applies transformations to the representation to address variations in the SQL dialect between the source and target systems, and finally converts the relational algebra representation back to the SQL dialect of the target system.
The Transport framework provides interfaces for common data types supported by each engine (e.g. $int$, $float$, $double$, $boolean$, $string$, $struct$, $array$, and $map$). We provide engine-specific implementations of Transport’s interfaces as wrappers over native types supported by the engine and define masking UDFs using Transport’s type system. The Transport framework for authoring UDFs provides multiple benefits:
  • Saves development time as developers need to author the UDF only once.
  • Simplifies code maintenance as we have a single source of truth for the UDF logic.
  • Ensures UDFs remain performant as Transport accesses native data types directly without needing conversion between different type systems.
6. Performance Evaluation
In this section, we evaluate the performance of the UDF implementing the data-masking operation. Specifically, we study the impact of factors such as: (1) data type of a column being masked, and (2) size and depth of the column being masked and (3) the number of columns being masked on the overall performance of the data-masking view.
6.1. Methodology
We first set up a micro-benchmark to evaluate the performance of the $MASK\_FIELD\_IF()$ UDF in a controlled and isolated environment. We analyze trends in the UDF’s performance as a function of the above-mentioned factors and provide insights obtained through CPU profiles. We confirm the trends observed from the micro-benchmarks with results obtained by running Spark SQL queries that use the $MASK\_FIELD\_IF()$ UDF on a large compute cluster running Apache Spark 3.1.1. Each query uses 50 executors, each provisioned with 8GB memory and 4 CPU cores.
We use a homegrown synthetic data generation tool to generate datasets that include the following types of columns common in our ecosystem:

### v2

@@ V-A1 Policy Compiler
A key step in the policy compiler is to convert the enforcement of a data access policy $p$ on a field path $f$ into a relational algebra plan representing the mask operator. We focus on the implementation of the mask operator to perform column-level masking i.e. the data element to be masked is inside an attribute of $\mathcal{R}$. A naive approach to implementing the data-masking operator for a relation-valued attribute is to unnest the attribute, mask the data at a given field path within the attribute and reconstruct the attribute into its original structure. However, this approach produces complex SQL queries for relation-valued attributes, which can be both hard to read and debug especially when involving transformations on deeply nested attributes. We therefore implement a scalar version of the mask operator with a user-defined function for column-level masking, which is maintained by Data Guard platform team.
Specifically, consider a field path $f$ inside an attribute $\mathcal{A}$ of a relation $\mathcal{R}$ and a boolean condition condcond, the following UDF implements a scalar version of the data masking operter maskmask: MASK_FIELD_IF(cond: BOOLEAN, $\mathcal{A}$: ANY, $f$: VARCHAR)
@@
If the condition is true, the UDF performs an in-place update on attribute $\mathcal{A}$ setting the value at path $f$ to NULL, instead of a sequence of unnest, transform, and nest operations. If the condition is false, the UDF returns the original value of $\mathcal{A}$. Therefore, the return type of the UDF is the same as the data type of $\mathcal{A}$ to ensure the view schema is identical to the underlying table schema.
The data-masking operator introduced in Equation 2 can thus be implemented via a simple UDF call in SQL as follows:
MASK_FIELD_IF(p.c, A, f)
where A is the root attribute of the field path $f$.
V-A2 Consent Bitmaps
@@
Recall from Section IV-A3 that the most common type of data access policies in our system are consent-based, where the policy condition involves checking preferences of data subjects stored in a separate table, referred to as the lookup table. As described in Section IV-B, a naive implementation of policy compiler can produce views that join a table containing data subject information with the lookup table using the data subject id as the join key. At LinkedIn, this table stores the consents of more than 1 billion members and any joins against this table result in expensive data shuffles. In Spark for instance, such joins are implemented as Sort Merge Joins since the size of both the tables involved in the join operation are too large to be broadcast using Broadcast Hash Joins. Optimizing the view SQL for fetching attribute values from the lookup table is therefore critical to minimize the cost of adopting views.
@@
In order to address this scalability challenge, we pre-compute a consent bitmap for every consent attribute referenced in a policy definition and leverage statistics associated with each consent to minimize the bitmap size. Specifically, we use a ”true bitmap” (bitmap of data subject identifiers whose consent value is true) if a minority of users have consent values set to true, and a false bitmap otherwise. Therefore, the number of user ids in each bitmap is at most half of the total number of users. Given that data subject identifiers at LinkedIn are numeric (a scenario we expect to be common), we use the Roaring Bitmap implementation  [37] to compute the bitmaps as it achieves superior compression ratios compared to other compression methods. As an example, a Roaring bitmap of 200 million 8-byte integers that represent user ids has a size of around 134MB, which is small enough for the bitmaps to be loaded into memory on the worker nodes in Spark and Trino. In cases where identifiers are non-numeric, we can employ minimal perfect hashing functions [38] to map keys into consecutive integers and then use Roaring Bitmap on the resulting integers.
@@
These bitmaps are then saved in a storage system like HDFS and loaded at access time for fast in-memory lookup via a UDF. The bitmap implementation is much cheaper compared to a join-based solution, as we only introduce a small memory overhead for storing bitmaps in each worker’s memory. In order to abstract away the storage system implementation details, we wrap the bitmap loading and lookup actions into a BitmapManager interface, thus ensuring that the bitmap solution can be adopted across a range of different storage systems.
@@
To ensure freshness of bitmaps, a bitmap computation job is scheduled periodically to reflect the latest user consents in the generated snapshot. The period is decided based on legal requirements. Each snapshot of a bitmap is addressed using a combination of the bitmap name and a timestamp when the bitmap was generated. The different snapshots of a bitmap are stored in a time-partitioned folder. The bitmap lookup is wrapped inside a UDF with the following signature: HAS_USER_CONSENT(consents: VARCHAR, user_id: BIGINT, access_time: TIMESTAMP), where:
@@
  • access_time: the timestamp for determining the specific snapshot of a bitmap to load. This parameter gives developers the ability to replay a query at a later time (aka time travel). By default, the data-masking views use the built-in SQL function CURRENT_TIMESTAMP() for the access_timeaccess\_time parameter. The access_timeaccess\_time parameter ensures that the same snapshot of bitmap is loaded across all the worker nodes for a given query.
V-A3 Masking deduplication
Given the scale of LinkedIn’s data warehouse and the diversity of policies that need to be enforced, it is common to find tables in our warehouse containing attributes that are subject to multiple policies. Further, a single policy may apply to multiple attributes within the table. A naive implementation of the policy compiler can result in redundant application of the MASK_FIELD_IF() UDF.
@@ V-A3 Masking deduplication
  • If a nested field needs to be masked based on one or more policy conditions, each of which is applicable to an ancestor, then a masking operation on the nested field is unnecessary. For example, masking operation on the field path in row 2 of Table I is redundant given an identical policy condition applicable to the root path \$. Similarly, the masking on field path in row 4 is redundant given an identical policy condition applicable to its parent in row 3.
  • If a field in a table has two or more applicable policies and the set of consents used in policy $p_{i}$ is a superset of the set of consents used in a different policy $p_{j}$, then the masking with policy $p_{j}$ is unnecessary. For example, rows 5 and 6 of Table I show two different policies $p_{3}$ and $p_{4}$ applied to the same field path with overlapping consents. It is easy to see that policy condition checks in row 2, 4, 6 are redundant and can be eliminated.
@@
[…]
2 $\mathdollar.col_{2}.field_{21}$ $consent_{1}$ $p_{1}$
3 $\mathdollar.col_{3}$ $consent_{2}$ $p_{2}$
4 $\mathdollar.col_{3}.[item].[?(@.field_{31}=^{\prime}s1^{\prime})]$ $consent_{2}$ $p_{2}$
5 $\mathdollar.col_{4}.[value]$ $consent_{3}$ AND $consent_{4}$ $p_{3}$
6 $\mathdollar.col_{4}.[value]$ $consent_{3}$ $p_{4}$
TABLE I: Maskings for different field paths with policies
To eliminate such redundant masking operations, we construct a schema tree from a given list of field paths and their applicable policies. A schema tree is one where each node represents an operator of a field path expression defined in Definition 2 and nodes corresponding to successive operators of a field path expression have a parent-child relationship between them. A node that is a final operator of a field path has an attribute to store the list of all policies applicable to that field path. Figure 5(a) shows an example schema tree constructed from field paths in Table I.
@@
Figure 5: A schema tree constructed from field paths in Table I.
After constructing the schema tree, we perform a pre-order traversal to prune redundant policies as detailed in Algorithm 1. The algorithm computes the minimal set of policies required to cover the set of consents of a node that do not overlap with any of its ancestors. This problem can be mapped to the well-known minimum set cover problem, which is an NP-complete problem [39]. We therefore use a greedy approximation algorithm that chooses policies with largest number of non-overlapping consents first. Figure 5(b) demonstrates the schema tree produced by applying Algorithm 1. Maskings in rows 1, 3, and 5 are chosen to be retained by the algorithm.
@@
    select $p\in P$ that maximizes $|consents(p)\cap\Delta|$
    $\Delta\leftarrow\Delta\setminus consents(p)$
@@
Suppose $col_{1}$ of $\mathcal{R}$ is labeled as a user id column, the resulting data-masking view SQL produced by the policy compiler is as follows:
@@ V-A4 View Maintenance
      If a view update is deemed necessary, the system invokes the policy compiler to generate a new view definition, and
@@
      registers the updated view definition with the view catalog (a Hive-based [40] catalog in our ecosystem).
To ensure smooth rollout of views and enable rollbacks in case of errors, each data-masking view is versioned according to the semantic versioning scheme [41]. Data pipelines are routed to the latest view versions by default. We also provide developers with the option to pin their pipelines to specific versions of views and be notified when newer versions of views become available. In this mode, developers can test their pipelines against the new view versions before deploying the change to production. Older versions of views are garbage collected periodically to minimize resource consumption due to stale views.
V-B View Consumption
At LinkedIn, developers write data processing applications using Apache Spark and Trino as the underlying compute engines. To ensure ease-of-use of data-masking views in applications, we implemented two important optimizations: (i) dynamic table-to-view routing, and (ii) multi-engine support, each of which will be discussed in detail in this section.
V-B1 Dynamic Table-to-View Routing
A key design choice of our system was to ensure access control policies are automatically enforced during data accesses by redirecting table accesses to the appropriate data-masking views determined by the purpose of access. This minimizes overhead for existing developers to ensure their accesses remain compliant. Recall from Section III that the table to view routing at access time is accomplished using a component called ViewShift. At LinkedIn, we implement this component as a plugin inside existing engine catalog implementations. ViewShift intercepts loadTable() API calls to the catalog from applications and returns the corresponding data-masking view instead. The ViewShift plugin exposes the following API:
@@ V-B1 Dynamic Table-to-View Routing
where contextMap parameter is query context map that contains attributes such as the purpose of the access (obtained from the identity of the developer that submits the query), the optional pinned version of the data masking view, and the environment where the application is running.
@@
V-B2 Multi-engine support
Data Guard effectively supports diverse data processing engines (Spark, Trino) with differing SQL dialects and UDF systems - a common challenge in large organizations. To ensure consistent enforcement across these systems without duplication, we leverage two open-source technologies built in-house:
@@ V-B2 Multi-engine support
Coral [42], our SQL translation tool, converts data-masking views from a centralized Hive SQL dialect to engine-specific dialects at access time using Apache Calcite [43]. It parses SQL from source dialect into a relational algebra representation, applies transformations to address variations in SQL dialects, and converts the representation into the target SQL dialect.
Transport UDF [44] framework enables writing core UDF logic once, and making it reusable across multiple engines through light wrappers. By abstracting engine-specific implementations behind common interfaces, Transport eliminates redundant code while maintaining native performance by directly accessing each engine’s type system without data conversions.
VI Performance Evaluation
When data-masking views are used to enforce access control policies, it is important to ensure that the performance overhead incurred during data access is minimal. In this section, we first study the performance overhead of data-masking views through a micro-benchmark to understand how the view performance is impacted given the access control policies being applied. We then share some statistics on the usage of data-masking views in our production warehouses at LinkedIn.
VI-A Micro-benchmark
Since there has been no prior standard benchmark to evaluate the performance of access control policies, we develop our own micro-benchmark to understand the performance characteristics of the data-masking operation implemented using the MASK_FIELD_IF() UDF. The primary goal of this micro-benchmark is to isolate and measure the performance overhead introduced by the data-masking view compared to the baseline of accessing the original table directly without any masking. In the micro-benchmark, we ignore post data-access operations such as joins and aggregations, as these operations are application-specific and orthogonal to the data-masking operation.
We developed a synthetic data generation tool to produce datasets with a variety of schemas and data types commonly used in practice, including: